OpenStudy (anonymous):

6^(6n) is less than equal to 1/81

myininaya (myininaya):

\[6^{6n} \le 3^{-4}\] \[\ln6^{6n}=\ln3^{-4}\] \[6n*\ln6=-4*\ln3\] \[n=\frac{-4\ln3}{6\ln6}=\frac{-2\ln3}{3\ln6}\]

myininaya (myininaya):

oops forgot my inequality

OpenStudy (nikita2):

\[6^{6n}\le1/81\]

myininaya (myininaya):

\[n \le \frac{-2\ln3}{3\ln6}\]

OpenStudy (anonymous):

thats not an option i have

myininaya (myininaya):

you can use alnb=lnb^a property

OpenStudy (anonymous):

what?...

myininaya (myininaya):

\[n \le \frac{\ln3^{-2}}{\ln6^3}\] \[n \le \frac{\ln(\frac{1}{9})}{\ln216}\] \[n \le \frac{\ln1-\ln9}{\ln216}\] \[n \le \frac{-\ln9}{\ln216}\]